# Learning Physics From Video: Unsupervised Physical Parameter Estimation for Continuous Dynamical Systems

### v1

@@ 4 Experiments
Figure 3: (a) Latent space estimation of the dynamic variable $z$ for the three datasets. The blue line shows the ‘ground truth’ value $z_{\text{real}}$ of the simulated dynamics, the model was trained with the dynamics of the continuous line while the dotted line is the predictions of the test set. (b) Encoder output $z$ vs. $z_{\text{real}}$ for every frame in (a). We find that our unsupervised loss leads to a good agreement between the predicted and ground truth dynamics.
4.1 Fully controlled video datasets
We propose to find solutions to inverse problems that extend beyond simple video motion dynamics. Therefore, to evaluate three different continuous dynamical systems involving motion, intensity and scaling. We describe the details about dimensionality in Appendix. A.1.
Figure 4: Example frames from our datasets. Each row shows a different dataset, corresponding to a different continuous dynamical system, and each column a different time sample.
Motion systems can be understood as problems where pixels have a translation or rotation transformation; classic examples include pendulums, springs, and celestial movements. For the motion system, we simulated a pendulum since the majority of related literature uses it as an example. In this dataset, the state variable is the angle of the pendulum $z_{real}=\theta$ $\in[-90.0^{\circ},90.0^{\circ}]$.
Intensity dynamics consists of the change in grayscale pixel values. This problem can be seen in nature in voltage imaging of spiking neurons or photonic crystals. To create this dataset, we generated an irregular shape and assigned all pixels inside the corresponding intensity of our dynamical system. We normalized the intensity dynamics to be in the range $z_{\text{real}}\in[0.2,1.0]$.
Scale in a video refers to changes in the total number of pixels corresponding to our object of interest. It can be related to a growing bacteria population or liquid diffusion. For this dataset, we created a filled circle centred in the middle of the image, where the radius is proportional to the dynamic variable. However, the scaling transformation is not symmetric; while one half of the circle grows, the other half becomes smaller and vice versa. Here, we normalized the range of the radius dynamics between $z_{\text{real}}=r\;\in[-10,10]$.
State-of-the-art methods are tested on simulated datasets like ours [9, 18, 36, 37, 41, 42], where the image consists in a fixed black background, while the individual objects in the scene are in different channels of the image, from now on we refer to this representation as a "mask". The pendulum system is a typical case of study, but for scale and intensity, there were no baselines available. In Figure 4, we illustrate the behaviour of the three different datasets.
4.2 Latent Space Evaluation
First, we train our model using the three synthetic datasets and evaluate the dynamics $z$ estimated by the encoder $E_{\theta}(\cdot)$. We compare the estimated state variable $z$ to its ground truth value $z_{real}$, which was used to generate the data. Following Eq. 2, we consider second-order dynamics for the three datasets, where the evolution of the state variable $z$ follows a dampened oscillation.
In Figure 3, we show that the model is capable of estimating the dynamics $z$ for all three datasets. Although oscillatory dynamics can be challenging for neural networks, we find that our unsupervised loss fits the dynamical behaviour reasonably, with small deviations from the ground truth. Figure 3(b) depicts the encoder output $z$ against the ‘ground truth’ dynamics $z_{real}$ for every frame the train and test set (dots). For better visualization, a straight line (solid lines) was fitted to each dataset.
Importantly, our model output has complete physical interpretability since we directly use the differentiable, second-order ODE (Eq. 2) in the physics block during training. Due to these physics priors, the model is able to generalize at test time to frames at time steps unseen during training. The network’s extrapolation of $z$ to unseen future time steps is shown using dashed lines in Figure 3a.
We observe that the most accurate results are obtained using the ‘Intensity’ dataset; this stems from the fact that the temporal information is in the pixel intensities instead of location, which means that in the intensity dataset, information about the dynamics is less discretized than the other datasets. Specifically, using 8-bit integer values for pixel intensities, the input can assume $256$ different values. In contrast, with motion and scaling, the dynamics are discretized by the pixel locations. In particular, for a $(50\times 50)$ frame size, the discretization of the dynamic variable is increasingly more impactful, especially as the oscillation amplitude is decreased. This effect is seen clearly in the latent space dynamics of the ‘Scale’ dataset in Figure 3a (red line), which displays discrete jumps. The discretization error also disproportionally affects the parameter estimation in the ‘Scale’ dataset, specifically the prediction of the parameter $\gamma_{1}$, which is discussed in the next section.
Motion Intensity Scaling Expected
@@ 4.2 Latent Space Evaluation
$\gamma_{0}$ $3.943\pm 0.0083$ $3.887\pm 0.0348$ $4.055\pm 0.0261$ $4.0016$
$\gamma_{1}$ $0.1441\pm 0.007$ $0.0889\pm 0.0098$ $0.9102\pm 0.012$ 0.08
Table 1: Parameter estimation accuracy. The right-most column shows the expected values. We show the mean and standard deviation of each learnable parameter in the physics block after training. The values are obtained over 7 different runs with different initializations.
4.3 Parameter estimation Accuracy
@@
In this section, we present the accuracy of the model in learning the parameters $\gamma$ of the dynamical system equations. As in section 4.2, we consider second-order governing equations as defined in Eq. 1. We note that $\gamma_{0}$ is the frequency and $\gamma_{1}$ is the damping factor of the oscillations.
Table 1 shows the model’s accuracy to estimate $\gamma$. We used equation  2 to simulate the dynamics; details of the ground truth are given in Appendix. A.3. On the one hand, The results show the model is pretty accurate in calculating $\gamma_{0}$; this parameter defines the oscillation frequency of the system; in figure  3, we validate the model is accurate in finding it. On the other hand, $\gamma_{1}$ is the damping factor of the model; the three experiments present different performances. However, according to the discretization discussed in section 4.2, the inaccuracy in the motion and scale dataset is anticipated.
@@ 4.4 Robustness and Stability
While previous work can reliably generate frame reconstructions using physics blocks in latent space, they often lack an analysis of the parameter estimation.Thus, it is not known if the generated frames correctly use the latent space information in an interpretable manner. In fact, it is known that the models presented in baselines are sensitive to initialization and may fail to converge [15, 30]. In this section, we evaluate the robustness of our model against changes in parameter initializations.
@@
We initialize the learnable parameters $\gamma$ of Eq. 2 in the interval $[-10.0,10.0]$ over multiple runs. The ground truth values used to generate the synthetic datasets were $\gamma_{0}=4$ and $\gamma=0.08$. In Figure 5, we show the convergence of the parameter prediction during training with different initializations. As can be seen, the model converges close to the ground truth values for each dataset, independent of the initialization. The shaded area highlights that while the trajectories may initially diverge, the model is capable of finding the correct parameters consistently. Table 1 shows the final accuracy of this experiment.
$\gamma_{0}$ $\gamma_{1}$
Motion
Refer to caption Refer to caption
Intensity
Refer to caption Refer to caption
Scale
Refer to caption Refer to caption
$epoch\;\rightarrow$
@@
Figure 5: Robustness of the parameters estimation against different initializations. The rows indicate the different dynamical systems, while the columns are the parameters to estimate. The vertical axis corresponds to the parameter value, and the horizontal axis is the epoch. Blue lines show the value of the estimated parameter $\gamma_{i}$ over training epochs. Since convergence was relatively fast, the horizontal axis is on a logarithmic scale for visibility. The shading highlights the variance of the trajectories before convergence.
Number parameters Time epoch $(s)$ Uses Decoder Inputs Masks
PAIG [18] 5.27M 252.72
PAIG w/o U-Net 4.78M 80.56
NIRPI [15] 75.42K 0.11
Ours 4.19M 0.95
Table 2: Relevant differences between our model and the baselines.
4.5 Baseline comparison
@@
In this section, we compare against the baselines PAIG [18] and NIRPI [15] discussed in Sec. 2. We did not use our datasets in the comparison since baselines are not designed to handle our intensity and scale setting. Therefore, for a fair comparison, we use the dataset first proposed in [18] and reused in [15]. The details of the equation and visualizations are given in Appendix A.4.
@@
In Table. 2, we present a size comparison of the models along with the training time for one epoch when using the setup proposed by the authors. PAIG does not require object masks as input since it learns the segmentation via a U-Net in its pipeline. Other baselines, as well as our model for multiple objects, need masked inputs and, therefore, do not employ a segmentation block. We, therefore, also consider PAIG with masked input and without the U-Net block in our comparisons.
@@
Next, we empirically demonstrate the baseline’s sensitivity to initial conditions: We train each model twice in the experiment and initialize the estimated parameter $k$ with values $1.0$ and $10.0$, respectively. The expected value is $k=2$ as used in the baseline papers [18, 15]. In Figure 6, we observe that different initializations fail to converge to the correct value of $k$ for the baselines, while our model is consistent and converges to the desired value accurately. Besides, this experiment shows our model is suitable for two dimensional problems with multiple objects.
@@
Figure 6: Robustness of the parameter estimation compared to baselines. For a fair comparison, we use the synthetic dataset created originally by the authors of the baseline papers to evaluate their models (see Appendix A.4). We plot the trajectories of the estimated parameter $k$ during training with different initializations for our model (green) and for the two baseline models (red, blue). Dotted lines correspond to an initial value of $k=10.0$, and solid lines to $k=1.0$. Our model converges robustly to the ground truth value of $k=2.0$.
4.6 Real-world video evaluation
For the real use cases we presented two different systems: 1. An LED light video recording with the constant brightness change over time 7, similar to the intensity problems previously studied. 2. A pendulum recording is used to validate the model in a realistic version of the synthetic dataset. For training these models, no masks were needed in comparison to baselines  [18, 15].
For an LED recording with constant change of brightness over time 7 shows that the model performs accurately (model predictions in blue) when compared with the ground truth intensity values manually extracted for each frame (red). In Figure 8, we present a more realistic use case where we do not have access to the ground truth or precise manual annotations for each frame. Fortunately, the length of the string $L$ is known to be 120 cm (string length is not visible in the video). The task in this experiment is to learn the value of $L$. Quantitatively, in Figure 8a, the model can reliably estimate the value of the length parameter $L$. Besides, in the latent space Figure 8b, we can see that the model can accurately predict the natural damped oscillations of the pendulum.
Figure 7: Estimating dynamics from a real-world linear intensity change video. (a) The estimated dynamics from the model (Latent Space) are compared with a fitted line and the manually extracted ground truth; the plot shows three dots where example frames of the video inputs are shown in (b). (b) Examples of the video showing the original frames and the resized input used in the grey scale where no masks are needed. The plot in (a) shows that the latent space can accurately estimate the intensity, capturing the global behaviour over time and following the indicated dynamics on a real video.

### v2

@@ 5 Experiments
$\gamma_{0}$ $\gamma_{1}$
Motion $3.943\pm 0.008$ $0.144\pm 0.007$
Intensity $3.887\pm 0.035$ $0.089\pm 0.010$
Scale $4.055\pm 0.026$ $0.910\pm 0.012$
GT $4.0016$ $0.08$
Figure 4: (a) Latent space estimation of the dynamic variable $z$ for the three synthetic datasets. The blue line shows the ‘ground truth’ value $z_{\text{real}}$ of the simulated dynamics. The model was trained with the dynamics of the continuous line while the dashed lines show the ground truth (blue) and predictions (yellow, red, green) on the extrapolated test set. (b) Parameter estimation accuracy. Rows 1-3: mean $\pm$ standard deviation of each learnable parameter in the physics block after training, bottom row: ground truth (GT). The values are obtained over 7 different runs with different initializations. We observe good agreement between the predicted and ground truth dynamics.
5.1 Synthetic Video Datasets
We start in a fully-controlled setting with three synthetic datasets involving motion, intensity, and scale, visualized in Figure 5. We describe dataset details in the supplementary material. State-of-the-art methods for physical parameter estimation commonly use simulated datasets [10, 20, 42, 44, 50, 51], where objects appear in different colors on a black background. The pendulum system is a typical case of study, but for scale and intensity, there were no baselines available.
Motion. We used a pendulum model popular in literature [20, 17, 7]. In this dataset, the state variable is the angle of the pendulum $z_{real}=\theta$ $\in[-180.0^{\circ},180.0^{\circ}]$.
Intensity. We consider time-varying grayscale pixel intensity of an irregular shape. We normalized the intensity dynamics to be in the range $z_{\text{real}}\in[0.2,1.0]$.
Scale. We use a filled circle centered in the middle of the image, where the radius is proportional to the dynamic variable. However, the scaling transformation is not symmetric; while one half of the circle grows, the other half becomes smaller and vice versa. We normalized the range of the radius dynamics between $z_{\text{real}}=r\;\in[-10,10]$.
Figure 5: Example frames from the synthetic datasets. Each row shows a different dataset, corresponding to a different continuous dynamical system, and each column a different time sample.
Dynamics Representation in Latent Space
First, we train our model using the three synthetic datasets and evaluate the dynamics $z$ estimated by the encoder $E_{\theta}(\cdot)$. We compare $z$ to its ground truth value $z_{real}$, which was used to generate the data. Following Eq. 2, we consider second-order dynamics for all datasets, where the evolution of the state variable $z$ follows a dampened oscillation.
Figure 4 shows our model is capable of estimating the dynamics $z$ for all three datasets. Although understanding oscillations in data can be challenging for neural networks [49], our unsupervised loss fits the dynamical behaviour reasonably, with small deviations from the ground truth.
Importantly, our model has complete physical interpretability since we use the differentiable, second-order ODE (Eq. 2) in the physics block. Due to this physics prior, the model is able to generalize at test time to previously unseen time steps. The network’s extrapolation of $z$ to unseen future time steps is shown using dashed lines in Figure 4(a).
We observe that the most accurate results are obtained using the ‘Intensity’ dataset. We attribute this to discretization error: 8-bit inputs of pixel intensity can assume $256$ different values. In contrast, with motion and scale, the dynamics are discretized by the pixel locations. In particular, for a $(50\times 50)$ frame size, the discretization of the dynamic variable is increasingly more impactful as the oscillation amplitude is decreased. This effect is seen clearly in the latent space dynamics of the ‘Scale’ dataset in Figure 4a (red line), which displays discrete jumps. The discretization error also disproportionally affects the parameter estimation in the ‘Scale’ dataset, specifically the prediction of the parameter $\gamma_{1}$, which is discussed in the next section.
Parameter Estimation Accuracy
@@ 5.1 Synthetic Video Datasets
To evaluate our model’s accuracy in estimating parameters $\gamma$, we consider second-order equations defined in Eq. 2, where $\gamma_{0}$ is the frequency and $\gamma_{1}$ is the damping factor of the oscillations. (See supplement for details.)
Figure 4b shows the $\gamma$ values learned by our model. We find that the model can accurately estimate $\gamma_{0}$ for all systems. Figure 4a serves as validation that the model indeed fits the correct frequency. The varying accuracy for $\gamma_{1}$ is attributed to the discretization error discussed in Section 5.1, as the accuracy decrease correlates well with the increased discretization in the motion and scale datasets.
@@ 5.2 Robustness and Stability
While previous work can reliably generate frame reconstructions using physics blocks in latent space, they often lack an analysis of the parameter estimation. Thus, it is not known if the generated frames correctly use the latent space information in an interpretable manner. In fact, it is known that the baseline models are sensitive to initialization and may fail to converge [17, 34]. In this section, we evaluate the robustness of our model against changes in parameter initializations.
@@
We initialize the learnable parameters $\gamma$ in the interval $[-10.0,10.0]$ over 7 runs. Ground truth values were ${\gamma_{0}=4}$ and ${\gamma_{1}=0.08}$. Figure 6 shows the convergence of the parameter estimation during training with different initializations. While the trajectories may initially diverge, the trained model consistently converges to the ground truth values for each dataset, independent of the initialization. The relatively low standard deviations of the final $\gamma$ estimates in Figure 4b highlights the stability of our model.
@@
Figure 6: Robustness of the parameter estimation against different initializations. The rows indicate the different dynamical systems, while the columns are the parameters to estimate. The vertical axis corresponds to the parameter value, and the horizontal axis is the epoch. Blue lines show the value of the estimated parameter $\gamma_{i}$ over training epochs. Since convergence was relatively fast, the horizontal axis is on a logarithmic scale for visibility. The shading highlights the variance of the trajectories before convergence.
PAIG [20] PAIG w/o U-Net NIRPI [17] Ours
Parameters 5.27M 4.78M 75.42K 4.19M
Time per    Epoch $[s]$ 252.72 80.56 0.11 0.95
Decoder Yes Yes Yes No
Mask No Yes Yes Yes No
Table 1: Comparison of baseline models evaluated on the public dataset from [20]. Our model only needs a mask if there are multiple objects present in the video.
5.3 Baseline Comparison on Synthetic Data
@@
The baselines PAIG [20] and NIRPI [17] are not designed to handle the intensity and scale settings of our synthetic datasets. Therefore, for a fair comparison, we use the synthetic dataset first proposed in [20] and reused in [17]. It consists of two MNIST digits moving following the spring equation parametrized by the spring constant $k$; the dataset was created using $k=2$. Details of the governing equation and visualizations are given in the supplement. In this dataset the dynamics is given by the $(x,y)$ position of each digit, therefore the dimensionality of the latent space is $d=4$.
@@
Table 1 presents a size comparison of the models along with training time per epoch. PAIG [20] employs a U-Net to learn the object masks and does not require them as input. Other baselines, as well as our model for multiple objects, need masked inputs and, therefore, do not employ a segmentation block. We also consider PAIG [20] with masked input and without the U-Net block in our comparisons.
@@
We empirically demonstrate the models’ sensitivity to initial conditions, we consider PAIG [20] as proposed by the authors with the U-NET. We train each model twice and initialize the estimated parameter $k$ with values $1.0$ and $10.0$, respectively. The expected value is $k=2$ [20, 17]. In Figure 7, we observe that different initializations fail to converge to the correct value of $k$ for the baselines, while our model is consistent and converges to the desired value accurately. This experiment demonstrates that our model can also successfully tackle problems with multiple objects.
@@
Figure 7: Robustness of the parameter estimation compared to baselines. For a fair comparison, we use the synthetic dataset created originally by the authors of the baseline papers to evaluate their models [20]. We plot the trajectories of the estimated parameter $k$ during training with different initializations for our model (green) and for the two baseline models (red, blue). Dotted lines correspond to an initial value of $k=10.0$, and solid lines to $k=1.0$. Our model converges robustly to the ground truth value of $k=2.0$.
Pendulum Torricelli Sliding block
$L\left[m\right]$ $k\left[\frac{\sqrt{m}}{s^{2}}\right]$ $a\left[\frac{m}{s^{2}}\right]$
PAIG $1.01\pm 0.03$ $1.01\pm 0.04$ $1.01\pm 0.04$ $0.99\pm 0.01$ $0.99\pm 0.01$ $0.97\pm 0.02$ $0.35\pm 0.03$ $0.38\pm 0.02$ $0.37\pm 0.04$
NIRPI $0.77\pm 0.33$ $0.84\pm 0.53$ $0.63\pm 0.38$ $0.21\pm 0.03$ $0.14\pm 0.04$ $0.16\pm 0.01$ $-0.09\pm 0.88$ $-0.01\pm 0.5$ $-0.06\pm 0.01$
Ours $0.51\pm 0.01$ $1.07\pm 0.2$ $1.30\pm 0.02$ $0.0094\pm 4e^{-4}$ $0.0132\pm 5e^{-4}$ $0.0167\pm 4e^{-4}$ $1.29\pm 0.1$ $2.70\pm 0.09$ $3.44\pm 0.19$
GT $0.45$ $0.90$ $1.50$ $0.0095$ $0.0128$ $0.0162$ $1.441$ $2.300$ $3.141$
LED Free fall scale
$\gamma$ $a\left[\frac{m}{s^{2}}\right]$
Small ball Medium ball Large ball
Ours $2.24\pm 0.36$ $0.97\pm 0.04$ $0.41\pm 0.04$ $15.0\pm 2.1$ $9.51\pm 1.27$ $10.22\pm 1.21$
GT $2.3$ $0.92$ $0.46$ $9.8$ $9.8$ $9.8$
Table 2: Parameter estimation accuracy on real-world videos. The table shows how closely each model estimates the ground truth (GT) parameter values, with mean and standard deviation calculated over five videos per setup. On (a) we show the so-called motion problems and compare against the baselines. On (b) since there are no baseline methods for non-motion scenarios, and thus, we could not evaluate baselines for the LED and free-fall scale videos. Highlighted values indicate the estimates that are closest to the expected values.
5.4 Real-world video experiments on Delfys75
We evaluate our model and the baselines on Delfys75 dataset and dynamics described in Section 4. Our model was trained without masks, while PAIG [20] needs to learn them, and NIRPI [17] requires them as input. For both baselines, we used the hyperparameters given in their respective papers. We trained our model for 500 epochs to ensure parameters converged for all experiments; the learning rate was chosen as described in Section 3. Further training details and an analysis of the latent space of our model are given in the supplement. Our model uses a prior $Q\sim\mathcal{N}(0,1)$ for $\mathcal{L}_{2}$ (Eq. 12), except for the Torricelli experiment, where the governing equation (Eq. 15) includes a square root that introduces conflict if $z<0$. To avoid this, we change the expected prior to $Q\sim\mathcal{N}(1,0.2)$. As explained in Section 3, this change leads to a different renormalization factor.
Each model was tested on five videos to compute the mean and standard deviation for each setting. For the pendulum and the LED, the parameter of each setting is estimated. For the sliding block, multiple values of the parameters result in the same dynamics; hence only the total acceleration $a=g(\sin(\alpha)-\mu\cos(\alpha))$ is estimated. The gravitational constant is also predicted in the dropped ball experiment. Baselines are excluded from intensity and scale experiments as they only support motion dynamics. Table 2 shows the ground truth (GT) and the learned value of each parameter.
On all settings, our proposed model performs parameter estimation relatively accurately compared to baselines. In particular, we observe that PAIG [20] estimates parameters only close to the initial value (1.0). We train all models on a single video at a time on each setting. This leads to a small training set compared to the 10,000 videos used in [20], illustrating the data efficiency of our model compared to PAIG [20]. On the other hand, parameter estimates for NIRPI [17] converge to varying values, but are on average lower accuracy than those fitted by the proposed model.
6 Discussion and Limitations
We present a novel model and dataset for unsupervised physical parameter estimation from video. While previous methods do not study phenomena other than motion, we go beyond motion and include a variety of dynamical systems.
The proposed model avoids frame prediction, which is challenging, but useful for visualization. However, decoder-based methods need to be designed for the particular dynamical system, as a transformation-agnostic decoder may struggle with high-quality reconstruction. Without decoders, our approach works across various dynamical systems.
In addition, we directly examine latent space predictions, unlike prior models [17, 20, 23], which focus on frame reconstruction and do not discuss the accuracy of predicted dynamics. Since parameter estimation is the main goal, reconstruction is simply a tool to define the unsupervised loss, and therefore, the latent space should be analysed closely.
Our model does not resolve the absolute scale of the state variable, for example, that the pendulum goes from $[-180.0^{\circ},180.0^{\circ}]$. Yet, thanks to the loss function, the model solves its own metric, ensuring assumptions made in Section 3. The choice of the prior target distribution simply normalizes the dynamics and should not affect performance, as seen in Section 5.4 with the Torricelli experiment. Baselines implicitly or explicitly do this normalization using the spatial transformer, forcing the prediction to be in the pixel metric. In contrast, all our dynamics depend on the prior distribution assumed by the KL-divergence.
Baseline methods are sensitive to $\gamma$ initialization. Our method may also converge to local minima with small learning rates and high initial parameters. Therefore, we enforce variance in the learning rate depending on the initial values of the parameters, which allow for wider search spaces.
Our model demonstrates good performance on real-world videos, successfully handling challenges such as shadows and perspective distortions, without the need for object masks. Additionally, our approach outperforms baseline models–which prioritize reconstruction tasks–on parameter learning. Using a spatial transformer, baselines rely on the positions and velocities predicted by the encoder, overlooking minor variations introduced by the governing equation.
Our new real-world dataset is versatile: While baseline models do not support non-motion dynamics and require object masks, they could still be trained on the motion-based subset of our proposed dataset with the accompanying masks. Thus, we believe our dataset will be helpful for evaluating future models on a common dataset, enabling more comprehensive comparisons in complex real-world settings.
Limitations. Our model is suited for continuous, autonomous differential equations. However, some systems, such as fluids, are described with more complex differential equations. In addition, we need to guarantee that Eq. 7 is differentiable. Taken together, our model needs to be extended to tackle more complex use cases, including multiple objects in a scene with independent dynamics.
Due to discretization, performance may vary with resolution. For example, small changes of the dynamic variable in the scale dataset were challenging to detect as the model cannot achieve sub-pixel accuracy.
We avoid predicting object masks in videos. Real-world experiments in Section 5.4 show that our model does not need masks when trained with complex backgrounds, whereas baseline methods still require them [20, 17]. However, when multiple objects interact, masks become necessary, as demonstrated in experiment 5.3. While learning masks is an effective solution [18, 20], it relies heavily on the spatial transformer limited to roto-translation problems.
Nevertheless, limitations are mostly related to the early stages of the problem, opening multiple opportunities for future works; with our work, we provide a base which can be extended to more complex problems in different scientific domains and initiate data sharing to evaluate future models.
Appendix A Appendix / Supplemental material
Figure 8: Dataset baseline. It shows the evolution of the spring dynamical system of two MNIST digits over a static CIFAR10 background. Figure adapted from [20].
A.1 Synthetic Dataset Dimensionality
Each experiment uses an image size of $(50\times 50)$ pixels, and the simulated dynamics is the same for the three datasets (Eq. 2) normalized with respect to the image size and maximum pixel intensity. Each dataset consists of 500 training samples with 20 frames, with a final numerical dimensionality of $(samples\times frames\times\#channels\times width\times height)=(500\times 20\times 1\times 50\times 50)$.
A.2 Training Hyperparameters
The experiments and baselines were executed on an NVIDIA 3080 GPU. For our model, implemented in PyTorch, the encoder was trained using the Adam optimizer [25] with a learning rate of $1\times e^{-2}$ and the Kaming weight initialization for MLP layers.
A.3 Simulation Details
Here we discuss the details of the dynamics simulation of the experiments in section 5.1. The equation  2 represents a harmonic oscillator with the closed solution:
$z(t)=Ae^{-\zeta t}cos(\omega t+\phi).$ (18)
$\gamma_{0}=\omega^{2}+\zeta^{2}=4.0016.$ (19)
$\gamma_{1}=2\zeta=0.08.$ (20)
A.4 Ablation study
We present an ablation study to show the effect of the KL-divergence (KLD) term in our loss function; we used the intensity experiment presented in the synthetic experiments section. Table 3 shows the comparison of learned parameters of the physical equations along with their expected, ground truth (GT) values. In addition, Figure 9 shows the convergence discussed in the methods section, where a shortcut for the model to optimize the mean squared error (MSE) in the latent space is to converge always to the mean value of the dynamic variable $z$.
Parameter MSE+KLD MSE GT
$\gamma_{0}$ 3.99 5.7 4
$\gamma_{1}$ 0.08 6.6 0.08
Table 3: Ablation comparison of the two-term losses. The model relaying only MSE cannot learn the expected values, while the KLD term allows it to get a proper estimation.
Figure 9: Comparison of the model trained with the MSE loss only (orange) and with our additional KLD loss term. The model relying just on MSE converges to the mean value of the dynamics, while our term guarantees that the model has the expected diversity.
A.5 Baseline Dataset
In this section, we describe the dataset used to compare the baselines. Both baselines were tested on the dataset published by [20] and also used in [17]. The equation of motion used for both systems is Eq. 21
$\vec{F}_{ij}=-k(\vec{p}_{i}-\vec{p}_{j})-l\frac{\vec{p}_{i}-\vec{p}_{j}}{|\vec{p}_{i}-\vec{p}_{j}|}.$ (21)
where $k$ is the spring constant and $l$ is the equilibrium distance.
A.6 Delfys parameter estimation
Below, we provide the details on how the physical parameters and measurement errors were obtained. All parameters with their errors are shown in Table 4.
Scenario Setting 1 Setting 2 Setting 3
Pendulum $L=45\pm 0.1$ [cm] $L=90\pm 0.1$ [cm] $L=150\pm 0.1$ [cm]
$\zeta=-0.059\pm 0.008$ $\zeta=-0.059\pm 0.008$ $\zeta=-0.059\pm 0.008$
Torricelli $k=0.0095\pm 10^{-4}$ $\left[\frac{\sqrt{m}}{s^{2}}\right]$ $k=0.0128\pm 10^{-4}$ $\left[\frac{\sqrt{m}}{s^{2}}\right]$ $k=0.0162\pm 2\cdot 10^{-4}$ $\left[\frac{\sqrt{m}}{s^{2}}\right]$
Sliding block $\alpha=20\pm 1$ [deg.] $\alpha=25\pm 1$ [deg.] $\alpha=30\pm 1$ [deg.]
$\mu=0.21\pm 0.04$ $\mu=0.21\pm 0.04$ $\mu=0.21\pm 0.04$
LED $\gamma=2.3$ $\gamma=0.92$ $\gamma=0.46$
Free fall $r_{0}=3.35\pm 0.1$ [cm] $r_{0}=6\pm 0.1$ [cm] $r_{0}=10\pm 0.1$ [cm]
$h_{0}=20\pm 0.1$ [cm] $h_{0}=20\pm 0.1$ [cm] $h_{0}=20\pm 0.1$ [cm]
$f=1451\pm 118$ $\left[\frac{\text{pixels}}{\text{m}}\right]$ $f=1451\pm 118$ $\left[\frac{\text{pixels}}{\text{m}}\right]$ $f=1451\pm 118$ $\left[\frac{\text{pixels}}{\text{m}}\right]$
Table 4: Physical parameters for each experiment and their measurement errors. The first parameter per scenario is changed between settings.
Pendulum. The two parameters to approximate are the length of the pendulum $L$ and the damping coefficient $\zeta$. The length $L$ can be trivially measured, and the estimation error is the spacing between marks on the used tape measure, in this case 0.1 cm. Assuming a small initial angle $\theta_{0}$, the horizontal offset $x$ of the pendulum can be described by:
$x=A\exp\left(-\frac{\zeta}{2}t\right)\cos(\alpha t-\phi).$ (22)
with $A$ the amplitude, $\alpha$ the frequency, and $\phi$ the initial phase. The peaks of the above curve then decay as ${x_{\text{peak}}=A\exp\left(-\frac{\zeta}{2}t\right)}$. The damping coefficient $\zeta$ can therefore be obtained by fitting linear regression to the function $\ln(x_{\text{peak}})$. This approximation is done on the setting with the longest length of the string, due to the lowest initial angle $\theta_{0}$. The error corresponds to the standard deviation over the 5 videos in this setting.
Torricelli. The only parameter to estimate in this scenario is $k$ relater to the water flow rate. Given the initial height of the water $h_{0}$, the final height of the water $h_{t}$, and the length of the video $t$, the parameter $k$ can be computed as:
$k=2\frac{\sqrt{h_{0}}-\sqrt{h_{t}}}{t}.$ (23)
The video clips were cut such that initial and final heights were correspondingly always $h_{0}=7$cm and $h_{t}=1$cm. The error estimate is obtained by computing the standard deviation of the parameter $k$ over the five videos for each setting.
Sliding block. The two parameters to be measured are the inclination angle $\alpha$ and the friction coefficient $\mu$. The inclination angle was set by varying the height of the top of the ramp. Investigating the recorded videos with a protractor showed that the angle was correct to within one degree. Using the sliding block equation as shown in the paper, the friction coefficient can be computed as:
$\mu=\tan(\alpha)-\frac{2s}{gt^{2}}.$ (24)
with $s=72.6$cm the total travel distance of the block and $t$ the duration of the video. Since the friction coefficient should be constant across all settings, the estimate and the error are computed over all 15 videos for this experiment. The error corresponds to the standard deviation.
LED. The only relevant parameter is the decay $\gamma$. This decay is controlled automatically, and thus, the value is exact.
Free fall. The radius $r_{0}$ is measured with a tape measure with an error of $0.1$cm. The initial distance from the camera $h_{0}=20$cm was measured likewise. The focal length if calculated as $f=\frac{r(0)}{r_{0}}h_{0}$ and the error is the standard deviation of the focal length over 15 recordings for this experiment. It should be noted that the units of the focal length are $\left[\frac{\text{pixels}}{\text{m}}\right]$ as the focal length makes the conversion between metric and pixel spaces. Also for this reason, the focal length cannot be taken from the producer’s spec sheet for the used camera.
A.7 Delfys75 real-video training details
For all videos, we adhered to the specifications outlined in the methods section. The model was trained over 500 epochs with an initial learning rate $lr=1e^{-2}$. Since each experimental group contained a different number of frames due to varying dynamics, the batch size and the number of input frames per sample were adjusted accordingly. The delta time ($dt$), defined as the time interval between frames, is determined by the camera’s recording speed in frames per second ($fps=60$). While the minimum possible ${dt=\frac{1}{fps}}$ was an option, it was not suitable for all experiments. When frame-to-frame differences were negligible ($x_{i}\approx x_{i+1}$), $dt$ was increased to ensure meaningful variations between frames for prediction.
Experiment Batch Size Frames per Sample $dt\>(s)$
Pendulum 64 20 $\frac{1}{10}$
Torricelli 64 20 $\frac{1}{10}$
Sliding Block 32 10 $\frac{1}{30}$
LED 32 20 $\frac{1}{60}$
Free Fall Scale all 4 $\frac{1}{30}$
Table 5: Hyperparameters used for training on Delfys75 experiments.